# CADmium: Fine-Tuning Code Language Models for Text-Driven Sequential CAD Design

### v2

@@ Abstract
Computer-aided design (CAD) is the digital construction of 2D and 3D objects, and is central to a wide range of engineering and manufacturing applications like automobile and aviation. Despite its importance, CAD modeling remains largely a time-intensive, manual task. Recent works have attempted to automate this process with small transformer-based models and handcrafted CAD sequence representations. However, there has been little effort to leverage the potential of large language models (LLMs) for sequential CAD design. In this work, we introduce a new large-scale dataset of more than 170k CAD models annotated with high-quality, human-like descriptions generated with our pipeline based on GPT-4.1. Using this dataset, we fine-tune powerful code-LLMs to generate CAD sequences represented in a JSON-based format from natural language descriptions, demonstrating the viability and effectiveness of this approach for text-conditioned CAD generation. Because simple metrics often fail to reflect the quality of generated objects, we introduce geometric and topological metrics based on sphericity, mean curvature, and Euler characteristic to provide richer structural insights. Our experiments and ablation studies on both synthetic and human-annotated data demonstrate that CADmium is able to automate CAD design, drastically speeding up the design of new objects. The dataset, code, and fine-tuned models are available online.
Data and Model: huggingface.co/collections/chandar-lab/cadmium-6866b402be81f39321af98d4
Code Repository: github.com/chandar-lab/cadmium
@@ 1 Introduction
To address these limitations, we introduce CADmium [^1] The name CADmium reflects the integration of “CAD” (Computer-Aided Design) and draws inspiration from the metal cadmium, known for its malleability and utility in precision applications., a novel approach that reformulates CAD generation as a purely text-to-text task. Our contributions are as follows. First, we develop a novel annotation pipeline using GPT-4.1 (OpenAI, 2025) that processes up to 10 multi-view images of 3D objects along with their construction sequences to generate annotations that combine natural language fluency with geometric precision. We apply our pipeline on 176,017 CAD sequences from DeepCAD (Wu et al., 2021) and release a new dataset with improved expert-quality annotations. Second, we fine-tune Qwen 2.5-Coder-14B (Yang et al., 2024), a state-of-the-art instruction-tuned code LLM, on our dataset to generate JSON-formatted CAD sequences from natural language prompts, leveraging the capabilities of the pre-trained model without the need for specialized embedding layers. Recent works have shown that fine-tuning LLMs can be effective for generating code used in visual content creation (Rodriguez et al., 2024, 2025), emphasizing their versatility in visual generative applications. We evaluate our approach against state-of-the-art CAD generation models, including Text2CAD (Khan et al., 2024), using both point-cloud and mesh-based metrics established in prior work. We also validate our models on the Fusion360 Reconstruction dataset (Willis et al., 2021) using annotations generated by our pipeline and the CADPrompt benchmark (Alrashedy et al., 2024) that contains 200 objects annotated by human experts. Third, we introduce novel metrics based on sphericity, mean curvature, Euler characteristic, and watertightness to more extensively evaluate the structural and topological characteristics of the generated objects. We highlight the human-like fluency and conciseness of CADmium’s expert-level text annotations relative to Text2CAD, supported by extensive empirical analysis. Furthermore, our fine-tuned LLM achieves competitive performance against the pre-trained Text2CAD model, with improvements across some evaluation metrics as model size increases.
@@ 2 Related Works
As a result, subsequent works shifted toward conditional generation from point clouds, images, or text inputs (Xu et al., 2024; You et al., 2024; Alam & Ahmed, 2024; Uy et al., 2022; Khan et al., 2024; Dupont et al., 2024). Progress in text-conditioned CAD generation was, however, hindered by the lack of datasets with textual annotations. As a solution, Text2CAD (Khan et al., 2024) proposed a novel annotation pipeline based on LLMs and vision-language models (VLMs) to process multi-view images and design histories from the DeepCAD dataset and generate annotations ranging from beginner to expert level. Text2CAD produced approximately 660,000 annotations covering all DeepCAD samples and is the only open-source dataset with geometrically precise annotations currently available for training text-conditioned CAD generation models. Nonetheless, beginner-level prompts often lack the details required to unambiguously reconstruct objects, and expert-level prompts are drastically verbose, redundant, or sometimes superfluous. Furthermore, the OmniCAD dataset of machine-generated text annotations for CAD objects, proposed by Xu et al. (2024), contains only beginner-level prompts.
In addition to annotating the DeepCAD dataset, Text2CAD introduced a supervised generation pipeline using a frozen BERT encoder (Devlin et al., 2019), adaptive text layers, and a transformer decoder with cross-attention between text and CAD embeddings. CAD-MLLM (Xu et al., 2024) improved over this framework by incorporating LLMs and conditioning the generations on multiple modalities, namely point clouds, images, and text. CAD-MLLM also introduced a new annotation pipeline and additional CAD samples parsed from the ABC dataset. In parallel, FlexCAD (Zhang et al., 2024) proposed a hierarchical text representation of CAD sequences to make the generation more controllable. With CADmium, we primarily aim to advance state-of-the-art performance in the text-to-CAD task, leaving multimodal integration and enhanced controllability for future work.
@@ 3.1 Datasets and Annotation Pipeline
We address the quality concerns of Text2CAD annotations by re-annotating all 176,017 samples from the DeepCAD library using the same multimodal inputs as Text2CAD. These inputs include the minimal JSON metadata extracted from DeepCAD using Onshape’s FeatureScript (ons, ), which improves human readability by eliminating redundant information, and 10 multi-view images rendered with Blender’s Python module (Blender Online Community, 2025) corresponding to the top, bottom, and 8 sides. We leverage the multimodal capabilities of GPT-4.1 to generate concise, natural-sounding yet geometrically precise descriptions. Additionally, we annotate 8,307 CAD objects from the Fusion 360 Reconstruction Dataset (Willis et al., 2021) not found in the ABC corpus from which DeepCAD is derived, offering a robust measure of generalization. This comprehensive re-annotation effort across both datasets involved processing a total of over 600 million input and output tokens with GPT-4.1. Importantly, only the expert-level annotations in Text2CAD provide enough geometric detail to unambiguously reconstruct objects. While it may be useful to generate plausible objects from incomplete descriptions in some cases, evaluating them is challenging as multiple outputs could satisfy a partial description. As a result, we generate exclusively expert-level prompts and restrict our comparisons to the expert-level descriptions in Text2CAD.
@@
Figure 2: Comparative analysis of CADmium and Text2CAD expert-level annotations. (a) Vocabulary growth as a function of token count demonstrates that Text2CAD has a limited vocabulary compared to CADmium. (b) Human-likeness, clarity, visual faithfulness given image renders, and completeness against the minimal JSON as measured by Gemma-3 12B (Team et al., 2025) indicates that CADmium descriptions tend to produce more natural-sounding albeit challenging descriptions. (c-e) Distribution of word counts, unique words, and digit lengths within numerical expressions per annotation shows that CADmium descriptions are more concise and diverse.
@@
Looking first at corpus-level statistics, CADmium’s annotations are drastically more concise, most being between 100 and 200 words, compared to Text2CAD’s longer and often redundant annotations, often exceeding 200 words (Figure 2c). Despite being shorter, CADmium’s annotations achieve a greater lexical variety, with a higher ratio of unique words per annotation (Figure 2d). The vocabulary growth curve also shows that CADmium uses a much broader range of words across annotations, reaching over 18,000 unique words or twice as many as Text2CAD (Figure 2a). This richer vocabulary creates more diverse descriptions and reduces the risk of overfitting to a limited set of words. Additionally, CADmium’s annotations represent numbers in a more natural, human-like way compared to the excessive precision of Text2CAD (Figure 2e). A side-by-side comparison between CADmium and Text2CAD descriptions is provided in Appendix A.
@@
While statistics provide a general feel for the quality of the corpus, they may not tell how individual descriptions compare head-to-head. Therefore, we conducted an LLM-as-a-judge evaluation using Gemma-3 12B (Team et al., 2025) on 10,000 annotation pairs (one from CADmium, one from Text2CAD). We looked at four key factors: human-likeness, clarity and readability, visual faithfulness to image renders, and completeness with respect to the minimal JSON description. CADmium was preferred for human-likeness and clarity while achieving comparable visual faithfulness. Text2CAD scored higher on completeness, which aligns with our design choice to prioritize natural-sounding concise annotations over the exhaustive, JSON-like detail found in Text2CAD. We argue that such verbosity, while improving completeness, is less representative of how humans naturally describe CAD models and may artificially simplify the task for language models. Details on the evaluation setup and prompts can be found in Appendix C.
@@ 5 Discussion and Conclusion
With the absence of large and good-quality human-annotated textual descriptions of CAD objects that are publicly available, it is necessary to utilize powerful and multimodal language models that can generate human-level annotations given the images and design history of CAD objects. In this regard our main goals were to further the state-of-the-art in terms of the quality of the text annotations, improve the training process by using instruction-tuned Code Qwen models, and augment the evaluation pipeline by including additional metrics that can provide deeper structural and topological insights (Appendix D). We also highlight some important limitations in our work. Although we demonstrated that GPT-generated prompts exhibit human-like characteristics, further investigation is needed to evaluate their generalization to diverse types of human prompts. Additionally, we observed a few outlier reconstructions with high Chamfer distances from the ground truths, which may be mitigated by increasing the training data and model size. In future work, we plan to extend our approach with a multimodal framework that facilitates editing CAD designs.
@@ Appendix C Prompt Details
To evaluate the quality of our CADmium annotations against Text2CAD expert annotations, we employed an LLM-as-a-judge setup using Gemma-3 12B [Team et al., 2025], as described in Section 3.1 and with results presented in Figure 2b. In this evaluation, 10,000 paired annotations (one from CADmium, one from Text2CAD expert for the same CAD model) were assessed. To mitigate positional bias, the CADmium and Text2CAD annotations were randomly assigned to the labels “1.” and “2.” within the user message presented to Gemma-3 12B. For optimizing the classification process and minimizing output variability, the model was explicitly instructed to respond with only a single digit (0, 1 or 2), allowing the judgment to be captured efficiently. The specific prompts used to instruct Gemma-3 12B for each of the four evaluation criteria are detailed as follows:
@@
Figure 11: Prompt provided to Gemma-3 12B for the LLM-as-a-judge evaluation, tasking it to assess and compare the human-likeness of two provided CAD model descriptions (one from CADmium, one from Text2CAD expert).
Figure 12: Prompt provided to Gemma-3 12B for the LLM-as-a-judge evaluation, tasking it to assess and compare the completeness of two provided CAD model descriptions against the ground truth minimal JSON representation of the corresponding 3D object.
Figure 13: Prompt provided to Gemma-3 12B for the LLM-as-a-judge evaluation, tasking it to assess and compare the visual faithfulness of two provided CAD model descriptions against multi-view image renders of the corresponding 3D object.
Figure 14: Prompt provided to Gemma-3 12B for the LLM-as-a-judge evaluation, tasking it to assess and compare the clarity and readability of two provided CAD model descriptions (one from CADmium, one from Text2CAD expert).

### v3

# \$*\$\$*\$footnotetext: Equal Contribution

@@ Abstract
Computer-aided design (CAD) is the digital construction of 2D and 3D objects, and is central to a wide range of engineering and manufacturing applications like automobile and aviation. Despite its importance, CAD modeling remains largely a time-intensive, manual task. Recent works have attempted to automate this process with small transformer-based models and handcrafted CAD sequence representations. However, there has been little effort to leverage the potential of large language models (LLMs) for sequential CAD design. In this work, we introduce a new large-scale dataset of more than 170k CAD models annotated with high-quality, human-like descriptions generated with our pipeline based on GPT-4.1. Using this dataset, we fine-tune powerful code-LLMs to generate CAD sequences represented in a JSON-based format from natural language descriptions, demonstrating the viability and effectiveness of this approach for text-conditioned CAD generation. Because simple metrics often fail to reflect the quality of generated objects, we introduce geometric and topological metrics based on sphericity, mean curvature, and Euler characteristic to provide richer structural insights. Our experiments and ablation studies on both synthetic and human-annotated data demonstrate that CADmium is able to automate CAD design, drastically speeding up the design of new objects. The dataset, code, and fine-tuned models are available online [^1] https://github.com/chandar-lab/CADmium [^2] https://huggingface.co/collections/chandar-lab/cadmium.
@@ 1 Introduction
To address these limitations, we introduce CADmium [^3] The name CADmium reflects the integration of “CAD” (Computer-Aided Design) and draws inspiration from the metal cadmium, known for its malleability and utility in precision applications., a novel approach that reformulates CAD generation as a purely text-to-text task. Our contributions are as follows. First, we develop a novel annotation pipeline using GPT-4.1 (OpenAI, 2025) that processes up to 10 multi-view images of 3D objects along with their construction sequences to generate annotations that combine natural language fluency with geometric precision, improving upon prior approaches like Text2CAD (Khan et al., 2024) by leveraging a single, more capable multimodal foundation model. We apply our pipeline on 176,017 CAD sequences from DeepCAD (Wu et al., 2021) and release a new dataset with improved expert-quality annotations. Second, we fine-tune Qwen 2.5-Coder-14B (Yang et al., 2024), a state-of-the-art instruction-tuned code LLM, on our dataset to generate JSON-formatted CAD sequences from natural language prompts, leveraging the capabilities of the pre-trained model without the need for specialized embedding layers. We acknowledge that the competitive performance achieved is significantly derived from the impressive pre-trained qualities of the Qwen 2.5-Coder base model, upon which our SFT efforts provide domain-specific refinement. Recent works have shown that fine-tuning LLMs can be effective for generating code used in visual content creation (Rodriguez et al., 2024; 2025), emphasizing their versatility in visual generative applications. We evaluate our approach against state-of-the-art CAD generation models, including Text2CAD (Khan et al., 2024), using both point-cloud and mesh-based metrics established in prior work. We also validate our models on the Fusion360 Reconstruction dataset (Willis et al., 2021) using annotations generated by our pipeline and the CADPrompt benchmark (Alrashedy et al., 2024) that contains 200 objects annotated by human experts. Third, we introduce novel metrics based on sphericity, mean curvature, Euler characteristic, and watertightness to more extensively evaluate the structural and topological characteristics of the generated objects. We highlight the human-like fluency and conciseness of CADmium’s expert-level text annotations relative to Text2CAD, supported by extensive empirical analysis. Furthermore, our fine-tuned LLM achieves competitive performance against the pre-trained Text2CAD model, with improvements across some evaluation metrics as model size increases.
@@ 2 Related Works
As a result, subsequent works shifted toward conditional generation from point clouds, images, or text inputs (Xu et al., 2024; You et al., 2024; Alam & Ahmed, 2024; Uy et al., 2022; Khan et al., 2024; Dupont et al., 2024). Progress in text-conditioned CAD generation was, however, hindered by the lack of datasets with textual annotations. The first large-scale effort toward textual supervision for CAD generation was Text2CAD (Khan et al., 2024), which introduced an LLM/VLM-based annotation pipeline over DeepCAD and produced 660k beginner- to expert-level descriptions. Although foundational, its beginner-level prompts are often high-level and underspecified, while the expert-level ones tend to be overly long, redundant, or stylistically unnatural (some examples are given in Appendix A). Subsequent works expanded this direction with increasingly complex pipelines. OmniCAD (Xu et al., 2024) generated machine-produced descriptions using InternVL2-26B. CADFusion (Wang et al., 2025) also released a text-to-CAD dataset which relies on GPPT-4o and human refinement. More recently, CADLLM/TCADGen (Liao et al., 2025) proposed a semi-automated human-in-the-loop pipeline that separately processes multi-view images, point clouds, and CAD parameters to obtain detailed micro- and macro-level descriptions, producing a richer dataset of expert-style annotations. Across these efforts, existing datasets diverge into two categories: (i) expert-level, geometrically grounded descriptions (e.g., Text2CAD, CADLLM) that enable unambiguous reconstruction and form the basis of our comparisons, and (ii) higher-level, abstract, or stylistic prompts (e.g., OmniCAD, beginner prompts in Text2CAD and CADFusion) that are not directly comparable due to their limited geometric specificity. We intend to generate high-quality CAD annotations that strike a balance: they must provide sufficient geometric detail to allow precise reconstruction, while remaining concise and plausibly human-written so as to reflect real-world design prompting scenarios.
In addition to annotating the DeepCAD dataset, Text2CAD introduced a supervised generation pipeline using a frozen BERT encoder (Devlin et al., 2019), adaptive text layers, and a transformer decoder with cross-attention between text and CAD embeddings. CAD-MLLM (Xu et al., 2024) improved over this framework by incorporating LLMs and conditioning the generations on multiple modalities, namely point clouds, images, and text. CAD-MLLM also introduced a new annotation pipeline and additional CAD samples parsed from the ABC dataset. In parallel, FlexCAD (Zhang et al., 2024) proposed a hierarchical text representation of CAD sequences to make the generation more controllable. With CADmium, we primarily aim to advance state-of-the-art performance in the text-to-CAD task, leaving multimodal integration and enhanced controllability for future work. Recent works have explored related but distinct avenues: CAD-Coder (Guan et al., 2025) focuses on image-to-CAD generation using a VLM to produce executable Python code, a different input modality and sequence representation than our text-to-text JSON-based approach. Similarly, CADFusion (Wang et al., 2025) explores distinct training strategies like combining sequential learning with visual feedback via DPO. Finally, Liao et al. (2025) proposed a dual-channel Transformer-based generator (TCADGen) alongside a refinement model (CADLLM) to convert textual design prompts into executable CAD modeling sequences.
@@ 3.1 Datasets and Annotation Pipeline
We address the quality concerns of Text2CAD annotations by re-annotating all 176,017 samples from the DeepCAD library using the same multimodal inputs as Text2CAD. These inputs include the minimal JSON metadata extracted from DeepCAD using Onshape’s FeatureScript (ons, ), which improves human readability by eliminating redundant information, and 10 multi-view images rendered with Blender’s Python module (Blender Online Community, 2025) corresponding to the top, bottom, and 8 sides. We leverage the multimodal capabilities of GPT-4.1 to generate concise, natural-sounding yet geometrically precise descriptions. Additionally, we annotate 8,307 CAD objects from the Fusion 360 Reconstruction Dataset (Willis et al., 2021) not found in the ABC corpus from which DeepCAD is derived, offering a robust measure of generalization. This comprehensive re-annotation effort across both datasets involved processing a total of over 600 million input and output tokens with GPT-4.1. Importantly, only the expert-level annotations in Text2CAD provide enough geometric detail to unambiguously reconstruct objects. While it may be useful to generate plausible objects from incomplete descriptions in some cases, evaluating them is challenging as multiple outputs could satisfy a partial description. As a result, we generate exclusively expert-level prompts and restrict our comparisons to the expert-level descriptions in Text2CAD, and CADLLM annotations.
@@
Figure 2: Comparative analysis of CADmium, CADLLM, and Text2CAD expert-level annotations. (a) Vocabulary growth as a function of token count shows that CADLLM Liao et al. (2025) produces substantially longer annotations, with a vocabulary expansion rate comparable to CADmium and larger than Text2CAD. (b) Human-likeness, clarity, visual faithfulness given image renders, and completeness against the minimal JSON as measured by Gemma-3 12B (Team et al., 2025) and Mistral Small 3.2 24B (Mistral, 2025) indicates that CADmium descriptions tend to produce more natural-sounding albeit challenging descriptions. (c-e) Distribution of word counts, unique words, and decimal places within numerical expressions per annotation shows that CADmium descriptions are more concise and diverse (e reports only CADmium and Text2CAD distributions, as CADLLM annotations contain exclusively integer values).
@@
Looking first at corpus-level statistics, CADmium’s annotations are drastically more concise, most being between 100 and 200 words, compared to Text2CAD’s and CADLLM’s longer and often redundant annotations, often exceeding 200 words (Figure  2c). Despite being shorter, CADmium’s annotations achieve a greater lexical variety, with a higher ratio of unique words per annotation (Figure  2d). The vocabulary growth curve also shows that CADmium uses a much broader range of words across annotations, reaching over 18,000 unique words or twice as many as Text2CAD, and comparably to CADLLM (Figure  2a). This richer vocabulary creates more diverse descriptions and reduces the risk of overfitting to a limited set of words. Additionally, CADmium’s annotations represent numbers in a more natural, human-like way compared to the excessive precision of Text2CAD, while CADLLM annotations contain exclusively integer values. (Figure  2e). A side-by-side comparison between CADmium and Text2CAD descriptions is provided in Appendix A.
@@
While statistics provide a general feel for the quality of the corpus, they may not tell how individual descriptions compare head-to-head. Therefore, we conducted an LLM-as-a-judge evaluation using Gemma-3 12B (Team et al., 2025) and Mistral Small 3.2 24B (Mistral, 2025) on 10,000 groups of three annotations (one from CADmium, one from Text2CAD, one from CADLLM Liao et al. (2025)). We looked at four key factors: human-likeness, clarity and readability, visual faithfulness to image renders, and completeness with respect to the minimal JSON description. CADmium was preferred for human-likeness and clarity while achieving comparable visual faithfulness. Text2CAD scored higher on completeness, which aligns with our design choice to prioritize natural-sounding concise annotations over the exhaustive, JSON-like detail found in Text2CAD. We argue that such verbosity, while improving completeness, is less representative of how humans naturally describe CAD models and may artificially simplify the task for language models. CADLLM annotations were typically rated lowest across all metrics, implying their descriptions are less readable and faithful. Notably, CADLLM explicitly separates appearance and parameter descriptions, whereas CADmium and Text2CAD integrate them into a single, interleaved description. The disparity in scores may therefore reflect the LLM judge’s preference for this integrated text format over a segmented approach. Details on the evaluation setup and prompts can be found in Appendix C.
@@ 4.3 Impact of Annotation Quality and Training/Evaluation Congruence
5 Limitations
While CADmium demonstrates the efficacy of fine-tuning code-LLMs for text-to-CAD generation, multiple limitations remain.
A primary limitation is the absence of large-scale human evaluation for the generated annotations. We employed an LLM-as-a-judge paradigm to scale our assessment, but this approach may carry intrinsic biases. LLM judges can exhibit preference leakage, favoring outputs that resemble their own training data or specific stylistic patterns, potentially preferring machine-generated text. To mitigate this evaluation bias, we include two families of LLMs, Gemma-3 12B and Mistral Small 3.2 24B, for judging and comparing our annotations with previous datasets.
Our dataset focuses on expert-level, geometrically precise descriptions to ensure unambiguous reconstruction. However, this may limit the model’s ability to generalize to abstract, vague, or purely functional descriptions often used by non-experts (e.g., “design a handle for a mug”). The model may struggle when instructions lack explicit geometric parameters. Further, the current scope is limited to sketch and extrude operations. Advanced CAD features such as chamfers and fillets operations are not currently supported, limiting the complexity of manufacturable objects. However, unlike the rigid structure of vectorized representations, text-based CAD sequences are inherently flexible, facilitating the addition of novel operations.
@@ 6 Discussion and Conclusion
With the absence of large and good-quality human-annotated textual descriptions of CAD objects that are publicly available, it is necessary to utilize powerful and multimodal language models that can generate human-level annotations given the images and design history of CAD objects. In this regard our main goals were to further the state-of-the-art in terms of the quality of the text annotations, improve the training process by using instruction-tuned Code Qwen models, and augment the evaluation pipeline by including additional metrics that can provide deeper structural and topological insights (Appendix D). We also highlight some important limitations in our work. Although we demonstrated that GPT-generated prompts exhibit human-like characteristics, further investigation is needed to evaluate their generalization to diverse types of human prompts. Additionally, we observed a few outlier reconstructions with high Chamfer distances from the ground truths, which may be mitigated by increasing the training data and model size. In future work, we plan to extend our approach with a multimodal framework that facilitates editing CAD designs. Overall, we wish to integrate some of the advancements from recent concurrent works, particularly those focused on multimodality and enhanced controllability with CADmium’s training methodology.
@@ Appendix C Prompt Details
To evaluate the quality of our CADmium annotations against CADLLM and Text2CAD expert annotations, we employed an LLM-as-a-judge setup using both Gemma-3 12B (Team et al., 2025) and Mistral 3.2 Small (Mistral, 2025), as described in Section 3.1 and with results presented in Figure 2b. In this evaluation, 10,000 paired annotations (one from CADmium, one for CADLLM, one from Text2CAD expert for the same CAD model) were assessed. To mitigate positional bias, the annotations were randomly assigned to the labels “1.”, “2.”, and “3.” within the user message presented to Gemma-3 12B. For optimizing the classification process and minimizing output variability, the model was explicitly instructed to respond with only a single digit (0, 1, 2, or 3), allowing the judgment to be captured efficiently. The specific prompts used to instruct LLMs for each of the four evaluation criteria are detailed as follows:
@@
Figure 11: Prompt for the LLM-as-a-judge evaluation, tasking the judge to assess and compare the human-likeness of two provided CAD model descriptions (one from CADmium, one from Text2CAD expert).
Figure 12: Prompt for the LLM-as-a-judge evaluation, tasking the judge to assess and compare the completeness of two provided CAD model descriptions against the ground truth minimal JSON representation of the corresponding 3D object.
Figure 13: Prompt for the LLM-as-a-judge evaluation, tasking the judge to assess and compare the visual faithfulness of two provided CAD model descriptions against multi-view image renders of the corresponding 3D object.
Figure 14: Prompt for the LLM-as-a-judge evaluation, tasking the judge to assess and compare the clarity and readability of two provided CAD model descriptions (one from CADmium, one from Text2CAD expert).